A 1.65x10-4 F capacitor is used in conjunction with a motor. How much

energy is stored in it when 119 V is supplied?


1
Expert's answer
2022-05-25T14:25:29-0400
"W=\frac{CV^2}{2}"

"W=\frac{1.65*10^{-4}*119^2}{2}=1.17\:\rm J"